OpenStudy (anonymous):

Express answer in exact form. Find the area of one segment formed by a square with sides of 6" inscribed in a circle. (Hint: use the ratio of 1:1:√2 to find the radius of the circle.

OpenStudy (anonymous):

\[diameter DB=\sqrt{6^{2}+6^{2}}=6\sqrt{2}\] \[radius r=3\sqrt{2}\] \[area of one segment =\frac{ \pi r ^{2} }{ 4 }-\frac{ 1 }{ 2 }*r*r=\frac{\left( \pi-2 \right)r ^{2} }{ 4 }\] plug the value of r and solve.

OpenStudy (anonymous):

my answer has to be in something divided by something pi- something

OpenStudy (phi):

surji gave you the answer

OpenStudy (anonymous):

how do I find r tho

OpenStudy (phi):

he told you. but you can figure it out yourself  how long is the diagonal ? you can use 45-45-90 triangles or you can use pythagoras a^2 + b^2 = c^2